OpenStudy (anonymous):

factor the polynomial f(x). Then solve the equation f(x) = 0. f(x)=x^3-3x^2-16x+48

myininaya (myininaya):

f(x)=x^2(x-3)-16(x-3) f(x)=(x-3)(x^2-16) f(x)=0 => (x-3)(x^2-16)=0 => x=3,-4,4